OpenStudy (anonymous):

Medal Find the LCM. 9, 8, 30

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ 9=3^2\hspace{.33em}\\~\\ 2=2^3\hspace{.33em}\\~\\ 30=2\times 3\times 5 \hspace{.33em}\\~\\ \text{take common multiples }\hspace{.33em}\\~\\ \text{lcm}=(2^3\times 3^2\times 5)=360 \hspace{.33em}\\~\\ }\end{align}\)